### v1

# Expanding RL with Verifiable Rewards Across Diverse Domains

Abstract
Reinforcement learning (RL) with verifiable rewards (RLVR) has shown promising results in mathematical reasoning and coding tasks where well-structured reference answers are available. However, its applicability to broader domains remains underexplored. In this work, we study the extension of RLVR to more diverse domains such as medicine, chemistry, psychology, and economics. We observe high agreement in binary judgments across different large language models (LLMs) when objective reference answers exist, which challenges the necessity of large-scale annotation for training domain-specific reward models. To address the limitations of binary rewards when handling unstructured reference answers, we further incorporate model-based soft scoring into RLVR to improve its flexibility. Our experiments show that a distilled generative reward model can serve as an effective cross-domain verifier, providing reliable reward signals for RL without requiring domain-specific annotations. By fine-tuning a base 7B model using various RL algorithms against our reward model, we obtain policies that outperform state-of-the-art open-source aligned LLMs such as Qwen2.5-72B-Instruct and DeepSeek-R1-Distill-Qwen-32B by a large margin, across domains in free-form answer settings. This also strengthens RLVR’s robustness and scalability, highlighting its potential for real-world applications with noisy or weak labels.
1 Instruction
Reinforcement learning with verifiable rewards (RLVR) has emerged as a promising training strategy for improving the reasoning capabilities of large language models (LLMs) (Luong et al., 2024; Lambert et al., 2024), even in the absence of a supervised fine-tuning stage (Guo et al., 2025). This is applied in a reference-based setting, where training assumes access to an objective ground-truth label for each prompt. A reward or verification function typically assigns a binary reward, indicating whether a model’s response aligns with the ground truth.
Previous research has primarily demonstrated the effectiveness of this method in tasks with well-structured answers in limited domains, such as mathematical reasoning and coding. In these settings, reference answers such as numerical values or predefined test cases enable straightforward rule-based verification. To improve the robustness of verifiers across tasks, recent studies have also explored training reward models within a single domain. For instance, researchers utilize approximately 800k instances to train a reward model for mathematical reasoning (Team et al., 2025).
In this work, we extend the study of RL with verifiable rewards beyond well-structured mathematics and coding tasks, investigating its potential to improve LLMs’ complex reasoning abilities across diverse domains, such as medicine, chemistry, psychology, economics, and education. Since training a separate reward model for each domain is impractical, we begin by querying an aligned, general-domain LLM, inspired by previous reference-free studies that demonstrated the superiority of generative verifiers (Zhang et al., 2024a; Mahan et al., 2024) over discriminative reward models in selecting the highest-scoring response from multiple samples during inference. Surprisingly, when provided with objective reference answers written by domain experts, we observe high agreement in binary judgments between closed-source and open-source LLMs (e.g., GPT-4o vs. Qwen2.5-72B-Instruct (Team, 2024)) across diverse domains. This high consistency suggests that, when high-quality, objective reference answers are available, reference-based grading is generally easier than reference-free assessment, which is inherently as difficult as identifying the first mistake in a response (Lightman et al., 2023). As a result, the necessity of large-scale annotation or synthesis of training data for single-domain reward models comes into question.
Furthermore, previous RLVR studies have predominantly relied on binary reward scores (Gandhi et al., 2024; Zhang et al., 2024b; Lambert et al., 2024; Guo et al., 2025; Ma et al., 2025a), raising concerns about their robustness in complex reasoning tasks across diverse domains. Even within mathematical problem-solving, our analysis of real-world examination problems shows that only $60.3\%$ of reference answers are single-word numerical values or expressions that can be relatively easily verified using rule-based reward functions. This proportion decreases further to $45.4\%$ for multi-domain problems. These observations underscore the limitations of binary scoring in assessing a broader range of reasoning-intensive tasks and highlight the potential benefits of exploring more expressive scoring mechanisms.
To address this, we incorporate model-based soft reward scores into RLVR by leveraging the probability of a single token that represents a generative verifier’s final judgment. Noticeably, we demonstrate that when objective reference answers are available, a reasonably effective verifier can be trained for diverse domains using a relatively small LLM, such as a 7B model, while achieving downstream performance comparable to much larger off-the-shelf generative verifiers, such as Qwen2.5-72B-Instruct (Team, 2024). In particular, this is achieved without the need for task-specific annotations for reward models. Instead, we collect policy’s responses and their assessments from a teacher verifier during online exploration. These judged responses, with varying performance levels and formatting noise, may further enhance the robustness of the learned reward model (Section 3.3). By fine-tuning a base 7B model using three RL algorithms guided by our reward model, we obtain policies that outperform powerful open-source aligned LLMs – such as Qwen2.5-72B-Instruct (Team, 2024) and DeepSeek-R1-Distill-Qwen-32B (Guo et al., 2025) – by up to $8.0\%$ in accuracy across various domains in free-form answer tasks. In addition to improving generalization, model-based rewards scale effectively with increasing data size. In contrast, rule-based rewards struggle, especially when addressing unstructured reference answers, and may even lead to performance degradation.
Our main contributions are as follows.
  • We extend RLVR to diverse domains, demonstrating its effectiveness beyond mathematics and coding, particularly in handling unstructured reference answers.
  • We show that even with unstructured objective answers, a relatively small cross-domain reward model can be trained without requiring task-specific annotations. [^1] Our reward model and free-form data will be released at https://huggingface.co/collections/virtuoussy/rlvr-67ea349b086e3511f86d1c1f.
  • We propose incorporating model-based soft rewards into RLVR and empirically demonstrate their superiority, scalability, and robustness over rule-based rewards across domains.
2 Related Work
2.1 Reward Estimation in Reinforcement Learning with Verifiable Rewards
For reasoning tasks such as mathematical reasoning, whether in constructing training data or at test time, a solution is typically considered correct if it arrives at correct final answer (Cobbe et al., 2021a). This is because reliably assessing the correctness of individual steps remains an open challenge, particularly when these steps may lack ground-truth labels in real-world scenarios. Similarly, the correctness of solutions to coding problems is typically accessed based on whether all test cases pass (Austin et al., 2021; Hendrycks et al., 2021a; Gehring et al., 2024). Consequently, previous reference-based RL studies have primarily focused on mathematical reasoning and coding tasks.
In most previous studies (Zelikman et al., 2022; Gandhi et al., 2024; Zhang et al., 2024b; Lambert et al., 2024; Guo et al., 2025; Ma et al., 2025a; Yu et al., 2025), given access to the reference answer $a$, the correctness label $z$ for a response $y$ to a prompt $x$ is typically a binary value. $z$ can also take on a value in the range $[0,1]$ to reflect varying degrees of correctness (Luong et al., 2024; Li et al., 2024; Ma et al., 2025b; Xie et al., 2025). Labels are assigned by a deterministic function $z=f(x,y,a)$, which operates based on predefined rules (e.g., exact match). These rules can also be combined with tools, such as a Python library, for verification (Xiong et al., 2025; Luo et al., 2025). This method is particularly effective when the answer type is fixed and easily matchable, such as a numerical value or a multiple-choice option. Each response is rated individually, without considering any preference information.
Besides using closed-source LLMs such as GPT-4o as verifiers (Chen et al., 2024), recent studies have also explored training reference-based reward models for mathematical reasoning (Team et al., 2025). However, these models are confined to a single domain and still require large-scale training data (e.g., $800$k instances for math) even within that domain.
2.2 Generative Reward Modeling
Using next-token prediction for reward modeling has attracted great interest in recent years (Lightman et al., 2023; Zheng et al., 2023; Tian et al., 2024; Zhang et al., 2024a), as it enables LLMs to fully leverage their generative capabilities, not only to produce accurate rewards but also to provide rationales that justify their judgments. In this work, we explore applying generative, reference-based verifiers to reinforcement learning and investigate their effectiveness across a variety of domains, an area that remains largely underexplored.
Furthermore, we explore training generative reward models without the need for annotated or synthetic step-by-step rationales (Team et al., 2025; Zhang et al., 2024a) to justify the final assessment. Specifically, we leverage the confidence of generative verifiers to provide stable and informative reward signals, enhancing the robustness of RL training in the presence of noise and ambiguity.
2.3 Verifiable Reasoning Data
Previous and on-going RLVR studies primarily focus on narrow tasks (Liu & Zhang, 2025; Xie et al., 2025) such as math word problem solving, code generation, and logic puzzles, where well-structured reference answers allow for straightforward rule-based verification. For example, SimpleRL (Zeng et al., 2025) and Tulu (Lambert et al., 2024) use math datasets GSM8K (Cobbe et al., 2021b) and MATH (Hendrycks et al., 2021b), in which each reference answer typically consists of fewer than two words. However, this reliance on well-structured data constrains the scale and diversity of resources that can be used for RLVR across broader domains.
In this work, we explore RLVR using reasoning data spanning diverse domains, where reference answers are free-form, either written by domain experts for unbiased evaluation (Yu et al., 2021), extracted from pre-training corpora (Yue et al., 2024), or generated by LLMs (Yuan et al., 2025).
3 Method
Figure 1: Overview paradigm of RLVR with our cross-domain verifier.
We focus on a setting where each prompt $x$ is accompanied by an expert-written reference answer $a$. Reference answers have been shown to play a crucial role in providing accurate rewards for reinforcement learning in reasoning-intensive tasks such as coding and mathematics (Shao et al., 2024). Ideally, in these domains, a response $y$ can be objectively verified against the given reference answer $a$. However, in practice, this verification process may be influenced by factors such as imperfect answer extraction and matching when pattern-based verifiers are used, as well as noise introduced by automated evaluation systems, such as a reward model $r_{\phi}(x,a,y)$.
Nevertheless, we can still use this verifiable reward in a policy gradient algorithm, with REINFORCE (Williams, 1992) as an example, as follows:
$J(\theta)=\operatorname{\mathbb{E}}_{(x,a)\sim D}\operatorname{\mathbb{E}}_{y_{i}\sim\pi_{\theta}(\cdot\mid x)}\big{[}r_{\phi}(x,a,y_{i})\Big{]}.$ (1)
When the generation of an entire response is modeled as a single action (Ahmadian et al., 2024), the gradient becomes (see Section A.3 for details):
$\nabla_{\theta}J(\theta)=\operatorname{\mathbb{E}}_{(x,a)\sim D}\operatorname{\mathbb{E}}_{y_{i}\sim\pi_{\theta}(\cdot\mid x)}\big{[}r_{\phi}(x,a,y_{i})\nabla_{\theta}\log\pi_{\theta}(y_{i}\mid x)\big{]}.$ (2)
3.1 Reward Estimation
To ensure a binary reward signal, we instruct a generative LLM $\pi_{\phi}$ to output only 00 or $1$ (see system prompt in Table A.1). For notational simplicity, we assume that each response consists of exactly $T$ steps, where each step corresponds to a non-empty line. Let $y_{i}^{T}$ denote the final step of response $y_{i}$. The binary model-based reward function is then defined as:
$r_{\phi}(x,a,y_{i})=\mathbbm{1}\big{(}c_{i}=1\big{)},$ (3)
where $c_{i}$ is sampled from $\pi_{\phi}(\cdot\mid x,a,y_{i}^{T})$, representing $\pi_{\phi}$’s judgment on the correctness of $y_{i}$.
Using $\pi_{\phi}$ as a verifier, we can also define a soft reward function using the probability of the judgment tokens (i.e., 00 or $1$):
$r_{\phi}(x,a,y_{i})=\begin{cases}\pi_{\phi}(1\mid x,a,y_{i}^{T})&\text{if }c_{i}=1,\\ 1-\pi_{\phi}(0\mid x,a,y_{i}^{T})&\text{if }c_{i}=0,\\ 0&\text{otherwise}.\end{cases}$ (4)
As shown in Equations 3 and 4, $r_{\phi}(x,a,y_{i})$ is bounded within $[0,1]$, ensuring consistency with the widely adopted binary reward scale.
3.2 Reward Normalization
To ensure stable gradients and encourage improvement across all samples in a batch that perform above average, we apply z-score normalization to rewards, inspired by prior studies such as GRPO (Shao et al., 2024) and REINFORCE++ (Hu, 2025).
$\tilde{r}(x,a,y_{i})=\frac{r(x,a,y_{i})-\mu_{r}}{\sigma_{r}},$ (5)
where $\mu_{r}$ and $\sigma_{r}$ denote the mean and standard deviation of the rewards within the batch containing $y_{i}$, respectively. In the special case where $\sigma_{r}=0$, we set all normalized rewards to zero, as these samples are either too difficult or too easy for the current policy.
3.3 Reward Model Training
When considering generative verifiers, a natural choice is to use an off-the-shelf aligned LLM as the reward model $\pi_{\phi}$, inspired by prior work that employs LLMs as judges (Zheng et al., 2023). However, we observe a noticeable performance gap on downstream tasks when using LLMs of different sizes. For example, the 72B reward model achieves 62.7% while the 7B model gets 58.8% on math data (see training details in Section 4). To address this, we explore training a moderately sized reward model (e.g., 7B) for general domains, aiming to balance performance and efficiency.
Since there are no ground-truth reward labels, for each $(x,a,y)$ triple, we prompt a fixed LLM to obtain the binary judgments $c\in\{0,1\}$, indicating whether $y$ matches the reference answer $a$. During the RL phase, we collect the data $\{(x,a,y,c)\}$ from the exploration stages and use it to fine-tune our reward models with supervised learning on $c$. Unlike relying on a fixed LLM to generate $y$, the improving actor policy produces responses with varying performance and potential formatting noise, which may enhance the robustness of the trained reward models.
4 Experiments
4.1 Data
Mathematics Data
To ensure high-quality reference answers, we use a large-scale dataset of 773k Chinese Question Answering (QA) pairs, collected under authorized licenses from educational websites. This dataset covers three educational levels: elementary, middle, and high school. Unlike well-structured yet small-scale benchmarks such as MATH (Hendrycks et al., 2021b) and GSM8K (Cobbe et al., 2021b), our reference answers are inherently free-form, often interwoven with rationales or involving several sub-questions yet lacking clear structural patterns. As a result, rule-based reward functions that rely on clean, well-structured answers for verification struggle to process these unstructured reference answers effectively.
We use GPT-4o-mini to translate questions and their corresponding responses into English. We randomly sample 3,000 QA pairs from each level and reserve them for testing. The average length of reference answers in the test set is 33.7, 36.3, and 53.9 words for elementary, middle, and high school levels, respectively. These are much longer than those in the GSM8K (1 word) and MATH (1.3 words) test sets.
Multi-Subject Data
Since no large-scale, free-form dataset with objective reference answers exists for general domains, we use a multi-subject multiple-choice QA dataset ExamQA (Yu et al., 2021). Originally written in Chinese, ExamQA covers at least 48 first-level subjects. We remove the distractors and convert each instance into a free-form QA pair. This dataset consists of 638k college-level instances, with both questions and objective answers written by domain experts for examination purposes. We also use GPT-4o-mini to translate questions and options into English.
For evaluation, we randomly sample 6,000 questions from ExamQA as the test set, while the remaining questions are used as the training pool. Since subject labels are not provided for each QA pair, we use GPT-4o-mini to classify them into one of 48 subjects or mark them as unclassified if uncertain. The detailed classification prompt is provided in Table A.1. Excluding unclassified instances (15.8% of the test data), the most frequent subjects include basic medicine, law, economics, management, civil engineering, mathematics, computer science and technology, psychology, and chemistry, as shown in Figure 2. For ease of analysis, we further categorize these subjects into four broad fields (STEM, social sciences, humanities, and applied sciences) as detailed in Table A.1. See examples in Table 10.
Figure 2: Distribution of subject occurrences in the test set of ExamQA (excluding unclassified).
Data for Training the Reward Model
We construct the data for training the reward model by extracting 20k samples from each training set of the two datasets, totaling 40k samples. Using the methodology in Section 3.3, we employ Qwen2.5-7B (Team, 2024) to conduct RL training. We use the RLOO (Kool et al., 2019; Ahmadian et al., 2024) algorithm and generate four online samples for each prompt. We use Qwen2.5-72B-Instruct as the reward model for hard label determination. By preserving all input-output pairs, this process yields 160k distilled training samples from Qwen2.5-72B-Instruct for reward model training.
To verify the training approach’s validity, we exclude these 40k original samples from the final training dataset. This strict separation ensures that the reward model never encounters any data used in previous training stages, thereby guaranteeing evaluation objectivity.
4.2 Baselines and Notations
Base  Directly use the base model to generate the response of the question.
SFT  Directly use the label (without CoT) to fine-tune the base model.
Rule-based reward  RL with the reward determined by predefined rules.
Qwen2.5-72B-Instruct  RL with the reward determined by the judgment of Qwen2.5-72B-Instruct (Team, 2024).
RM-7B (ours)  RL with the reward determined by the judgment of the reward model trained on our 160k distilled data based on Qwen2.5-7B-Instruct (Team, 2024).
Binary  When using rule-based rewards, we directly judge if the label is in the answer. When using model-based rewards, we use the output of the model. The value of binary reward should be in $\{0,1\}$.
Soft  When using rule-based rewards, we use Jaccard similarity  (Jaccard, 1912) as the reward. When using model-based rewards, we use the probability of the first output token. The value of soft reward should be in $[0,1]$.
4.3 Evaluation
We begin by investigating majority voting using a strong open-source LLM, Qwen2.5-72B-Instruct (Team, 2024), as the reward model $\pi_{\phi}$. The evaluation process follows the prompting template provided in Table A.1. Given a prompt $x$ and a reference answer $a$, we generate $m$ evaluation samples and determine the correctness of a response $y$ via majority voting. A response is considered correct if at least half of the evaluations classify it as such, i.e., $\sum_{j=1}^{m}\mathbbm{1}\big{[}\pi_{\phi}^{(j)}(x,y^{T},a)=1\big{]}\geq\frac{m}{2}$.
We measure the agreement between the Qwen-based evaluation method (majority voting over $m$ samples) and GPT-4o (a single evaluation per response) using Cohen’s Kappa ($\kappa$). As shown in Figure 3, the two evaluation methods demonstrate almost perfect agreement ($0.81\leq\kappa\leq 1.00$), with $\kappa$ exceeding $0.86$ for mathematics and $0.88$ for multi-subject college-level problems. This high level of agreement remains consistent across varying values of $m$, indicating that the results are not highly sensitive to the number of evaluation samples. Based on this observation, we adopt $m=1$ in all subsequent evaluations to improve efficiency without compromising evaluation quality.
4.4 Implementation Details
After obtaining the 160k distilled data from Qwen2.5-72B-Instruct, we perform supervised fine-tuning on Qwen2.5-7B-Instruct using this data, resulting in our reward model. We use different RL algorithms to validate the effectiveness of our method, including REINFORCE (Williams, 1992; Ahmadian et al., 2024), RLOO (Kool et al., 2019; Ahmadian et al., 2024), and REINFORCE++ (Hu, 2025). Following Stiennon et al. (2020); Ouyang et al. (2022); Hu (2025), we introduce a Kullback-Leibler (KL) divergence penalty between the RL model and the reference policy (i.e., base model) distributions to mitigate bias in the reward model. We update $\tilde{r}(x,a,y_{i})$ as follows:
$\tilde{r}(x,a,y_{i})\leftarrow\tilde{r}(x,a,y_{i})-\beta\log\Bigg{(}\frac{\pi_{\theta}(y_{i}\mid x)}{\pi_{\text{ref}}(y_{i}\mid x)}\Bigg{)},$ (6)
where $\beta\geq 0$ controls the effect of the KL penalty, and $\pi_{\text{ref}}$ represents the reference policy distribution. We set $\beta=0.01$ for all experiments.
For all algorithms, we apply reward normalization as introduced in Section 3.2. We use Qwen2.5-7B (Team, 2024) as the base model for our experiments. Despite not undergoing post-training, it demonstrates reasonable instruction-following capabilities, as shown by its zero-shot performance in Table 1. We also include the results of Qwen2.5-72B-Instruct and DeepSeek-R1-Distill-Qwen-32B to illustrate the difficulty level of our datasets. For both datasets, we select 30k samples as the training data. The training hyper-parameters of RL distilled data collection, reward model training, and the main experiments can be found in Table 9 in the Appendix.
Method Reward Score Type Math Multi-Subject
E M H Avg STEM Social Humanities Applied Others Avg
Qwen2.5-72B-Instruct 44.2 57.7 40.3 47.4 25.2 20.1 28.7 20.5 21.0 22.6
DeepSeek-R1-Distill-Qwen-32B 27.6 34.8 17.4 26.6 23.2 21.8 26.7 20.5 18.5 21.7
Base 43.1 53.9 33.2 43.4 16.3 14.9 15.2 13.3 14.8 15.0
SFT 53.6 50.5 32.9 45.7 24.6 22.8 25.7 20.9 22.6 23.1
REINFORCE rule based binary 58.5 66.5 46.7 57.2 25.3 26.6 27.7 21.1 20.7 24.2
soft 46.0 47.7 31.5 41.7 22.0 20.3 23.1 16.9 20.5 20.3
Qwen2.5-72BInstruct binary 64.4 72.1 51.6 62.7 27.9 27.9 30.7 24.4 23.2 26.6
soft 62.5 71.2 53.1 62.3 32.2 32.8 36.0 24.9 27.9 30.3
RM-7B (ours) binary 63.8 71.7 51.9 62.5 29.0 29.1 28.4 23.8 24.8 27.3
soft 62.9 70.7 53.0 62.2 32.7 32.8 35.6 28.6 27.4 31.2
REINFORCE++ rule based binary 56.4 65.5 47.6 56.5 26.1 26.1 26.4 21.8 24.7 25.0
soft 49.4 52.9 36.2 46.2 22.5 22.0 25.7 18.6 20.2 21.4
Qwen2.5-72BInstruct binary 63.0 71.3 50.4 61.6 30.7 32.8 34.3 27.5 27.8 30.3
soft 62.7 70.4 50.5 61.2 30.8 30.1 33.7 25.6 25.4 28.8
RM-7B (ours) binary 63.1 71.3 51.5 62.0 30.2 30.8 31.0 26.6 26.3 29.1
soft 62.7 70.3 50.8 61.3 29.5 31.7 33.7 25.8 26.2 29.0
RLOO rule based binary 58.2 67.0 50.2 58.5 28.2 27.9 27.4 22.4 24.5 26.3
soft 49.6 50.3 33.9 44.6 16.7 17.3 18.8 14.5 16.9 16.6
Qwen2.5-72BInstruct binary 63.0 70.8 51.1 61.6 29.4 30.5 33.7 24.6 26.1 28.4
soft 63.8 71.0 52.4 62.4 32.9 31.4 34.7 27.7 26.8 30.6
RM-7B (ours) binary 63.4 71.8 53.8 63.0 29.3 29.0 33.3 25.8 25.6 28.1
soft 63.3 71.7 53.6 62.9 31.0 32.0 35.6 27.0 27.0 30.0
Table 1: Performance Comparison of Different Methods. Base model: Qwen2.5-7B. E: elementary. M: middle. H: high.
4.5 Main Results
Table 1 shows the results on mathematics and multi-subject tasks. We have the following observations:
Evaluation on Base Models  Both our math and multi-subject data have demonstrated notable difficulty, with even strong open-source models like Qwen2.5-72B-Instruct (Team, 2024) and DeepSeek-R1-Distill-Qwen-32B (Guo et al., 2025) performing unsatisfactorily, particularly on multi-subject tasks (21.7% for DeepSeek-R1-Distill-Qwen-32B and 22.6% for Qwen2.5-72B-Instruct). We believe that more challenging datasets will better facilitate exploration across the industry.
SFT vs. RL  SFT significantly underperforms RL on both math and multi-subject tasks. Notably on math, SFT merely improves the model performance from 43.4% to 45.7%, falling far short of rule-based reward RL (RLOO, 58.8%) and lagging even further behind model-based reward RL (RM-7B, 63.0%). These findings demonstrate RL’s distinct advantages and potential in reasoning tasks when there is no high-quality Chain-of-Thoughts for training.
Model-based Reward vs. Rule-based Reward  From the table, we can conclude that model-based reward consistently outperforms rule-based reward in free-form reference-based scenarios. For instance, RM-7B (ours) and Qwen-2.5-72b-Instruct with binary reward achieves 63.0% and 61.6% respectively on average with RLOO, while rule-based reward only gets 58.5%. Notably, our distilled 7B reward model exhibits competitive performance against its much larger predecessor, Qwen2.5-72B-Instruct. In multi-subject evaluations using REINFORCE, the model trained from RM-7B achieves 31.2% accuracy compared to the 72B model’s 30.3% – a significant improvement given the substantial parameter disparity. This enhanced capability likely emerges from stabilized response patterns developed during training, which better align with the generative reward model’s objectives compared to the base model’s more variable outputs.
Binary Reward vs. Soft Reward  For rule-based reward, soft reward consistently underperforms binary reward. This discrepancy may stem from redundant tokens between the model’s generated answers and reference labels, which can lower the reward scores for correct answers. A potential improvement could involve adopting metrics like cosine similarity of sentence embeddings as soft rewards, as these may better capture semantic alignment. In contrast, for model-based reward, binary and soft rewards yield comparable results on math tasks. This suggests that the model likely produces judgments with extremely high confidence, as determining answer-label matches in mathematical problems is relatively easy. However, in multi-subject tasks, where reference labels exhibit greater diversity and consequently higher judgment complexity, soft rewards demonstrate more conservative scoring behavior. This conservatism in ambiguous cases enables soft rewards to outperform binary rewards in certain scenarios (31.2% vs. 27.3%, REINFORCE, RM-7B), as their soft scoring better accommodates the inherent uncertainty of open-domain evaluation.
Summary  Our method establishes new state-of-the-art performance in RLVR through three key innovations: (1) Our proposed model-based reward is much stronger than rule-based baseline, allowing various RL methods to obtain very accurate rewards in general domain scenarios. (2) Building upon the data distilled from Qwen2.5-72B-Instruct, we develop a computationally efficient 7B model that can achieve comparable or even better performance. (3) We extend binary reward to soft reward, which can get more conservative scores for ambiguous cases, which can help get better performance when the reference answers exhibit greater diversity and consequently higher judgment complexity.
4.6 Scaling Experiments
Method Scale Math Multi-Subject
E M H Avg STEM Social Humanities Applied Others Avg
Rule based 20k 58.9 68.1 47.6 58.2 27.3 28.0 31.4 23.5 23.0 26.2
40k 61.5 69.4 55.4 62.1 25.1 24.8 27.4 21.0 23.0 24.0
60k 62.6 69.8 56.8 63.1 20.0 21.9 26.4 16.6 19.9 20.1
80k 62.4 68.2 53.6 61.4 19.2 18.3 26.7 15.1 16.4 18.0
100k 52.6 57.2 45.2 51.7 17.8 18.2 20.5 13.4 16.4 16.9
RM-7B (ours) 20k 64.9 71.8 53.4 63.4 30.8 34.6 31.7 28.0 27.7 30.8
40k 65.6 72.4 54.4 64.1 34.3 33.7 36.3 29.5 28.6 32.4
60k 66.0 71.6 53.2 63.6 33.3 36.6 37.3 31.5 28.9 33.3
80k 66.6 72.3 55.6 64.8 34.5 38.6 38.3 31.6 31.0 34.6
100k 67.1 72.3 55.6 65.0 35.1 38.5 39.3 32.7 30.7 35.0
Table 2: The results of the scaling experiments. We use RLOO as the RL algorithm and binary reward as the score type.
Scalability has emerged as a critical property in the RL-based training era. A key question worthy of investigation is whether model performance can continue to improve as RL training progresses and data volume increases. To examine this, we conduct experiments using our trained reward model against rule-based reward while progressively scaling the dataset. We randomly sampled 100k samples from our training corpus as the baseline set, conducting evaluations on both mathematical reasoning and multi-subject tasks. Table 2 shows the experimental results.
The results reveal significant differences in scaling capabilities. The rule-based reward demonstrates unstable scalability across both mathematical and multi-subject tasks, exhibiting substantial performance fluctuations and eventual degradation as RL training continues. In contrast, our learned reward model shows consistent improvement trends throughout the training process. This empirical evidence highlights the inherent scalability advantages of model-based rewards compared to rule-based rewards.
4.7 Out-of-Distribution Evaluation
Method Natural Reasoning WebInstruct
Rule based 29.4 33.9
RM-7B (ours) 39.8 44.0
Table 3: The results of the Out-of-Distribution evaluation
To further validate the effectiveness of our reward model, we conduct additional evaluations on two benchmarks: NaturalReasoning (Yuan et al., 2025) and WebInstruct (Yue et al., 2024). We compare the performance of the rule-based reward with our RM-7B using RLOO with binary reward. The base model is Qwen2.5-7B. The results are shown in Table 3. As can be seen from the table, the performance of RM-7B remains significantly superior to the rule-based reward on datasets from other domains. This demonstrates that our general-purpose reward model can extend to other domains while maintaining strong performance.
5 Discussions and Conclusions
In this work, we simplify the verification task by instructing a generative reward model to output either $1$ or 00, without requiring chain-of-thought (CoT) reasoning (Nye et al., 2021; Wei et al., 2022). While CoT has proven useful in both reference-based (Team et al., 2025) and reference-free (Zhang et al., 2024a) settings, it remains an open question how necessary in-depth rationales are for assessing semantic equivalence between reference answers and model responses in the same language, particularly when focusing on the conclusive part of each response. This also raises a related question for process reward modeling (Lightman et al., 2023) in RLVR: how should rewards be assigned when there is no direct supervision for intermediate steps, regardless of the step segmentation method?
In addition, we do not consider format-based rewards (Guo et al., 2025) in this work. We revisit the role of format-related constraints and rewards in this context. In prior work, pattern-based functions are often used for scoring, making it critical to guide LLMs to enclose their final answers in an easily parsed format. These extracted answers are then compared with the reference answers for verification and evaluation. In contrast, by reintroducing a reward model in RLVR without imposing any format constraints on reference answers or model responses, we reduce the need for extensive human effort in data standardization and pattern design.
Appendix A Appendix
A.1 Template
Table A.1 shows the template for the grading task. Table A.1 shows the template for the classification task. Table A.1 shows the classification of subjects into STEM (Science, Technology, Engineering, and Mathematics), Social Sciences, Humanities, and Applied Sciences.
Given a problem, determine whether the final answer in the provided (incomplete) solution process matches the reference answer. The reference answer may be one single option character (e.g., A, B, C, D), a numerical value, an expression, or a list of answers if multiple questions are involved. **The reference answer may be in Chinese or another language, but your evaluation should be language-agnostic.** Your task: - Compare the final output of the solution process with the reference answer. - If they **match exactly**, output **YES**. - If they **do not match**, output **NO**. - If the solution process is unclear, incomplete, or ambiguous, assume it is incorrect and output **NO**. Your output must be strictly **’YES’** or **’NO’**, with no additional words, punctuation, or explanation. --- **Question:** {question} **Solution Process (Final Step Only):** {response} **Reference Answer:** {reference} **Output:**
Table 4: Template for the grading task.
Based on the content of Question and Answer classify the subject into one of the following categories. Return only the corresponding subject ID. If classification is uncertain, return 999. **Question:** {question} **Answer:** {answer} 110 Mathematics 120 Information Science and System Science 130 Mechanics 140 Physics 150 Chemistry 170 Earth Science 180 Biology 190 Psychology 210 Agronomy 230 Animal Husbandry and Veterinary Science 310 Basic Medicine 320 Clinical Medicine 330 Preventive Medicine and Public Health 350 Pharmacy 360 Chinese Medicine and Chinese Materia Medica 413 Information and System Science Related Engineering and Technology 416 Natural Science Related Engineering and Technology 420 Surveying and Mapping Science and Technology 430 Materials Science 460 Mechanical Engineering 470 Power and Electrical Engineering 510 Electronics and Communications Technology 520 Computer Science and Technology 530 Chemical Engineering 550 Food Science and Technology 560 Civil Engineering 570 Water Conservancy Engineering 580 Transportation Engineering 610 Environmental/Resource Science and Technology 620 Safety Science and Technology 630 Management 710 Marxism 720 Philosophy 730 Religious Studies 740 Linguistics 750 Literature 760 Art 770 History 790 Economics 810 Political Science 820 Law 840 Sociology 850 Ethnology and Cultural Studies 860 Journalism and Communication 870 Library, Information, and Documentation 880 Education 890 Sports Science 910 Statistics 999 Unclassified
Table 5: Template for the classification task, with subject names and IDs referenced from (Yu et al., 2021).
Category Subject IDs
STEM 110 (Mathematics), 120 (Information Science and System Science), 130 (Mechanics), 140 (Physics), 150 (Chemistry), 170 (Earth Science), 180 (Biology), 430 (Materials Science), 460 (Mechanical Engineering), 470 (Power and Electrical Engineering), 510 (Electronics and Communications Technology), 520 (Computer Science and Technology), 530 (Chemical Engineering), 560 (Civil Engineering), 570 (Water Conservancy Engineering), 580 (Transportation Engineering), 610 (Environmental/Resource Science and Technology), 620 (Safety Science and Technology), 910 (Statistics)
Social Sciences 190 (Psychology), 790 (Economics), 810 (Political Science), 820 (Law), 840 (Sociology), 850 (Ethnology and Cultural Studies), 860 (Journalism and Communication), 870 (Library, Information, and Documentation), 880 (Education), 890 (Sports Science), 630 (Management)
Humanities 710 (Marxism), 720 (Philosophy), 730 (Religious Studies), 740 (Linguistics), 750 (Literature), 760 (Art), 770 (History)
Applied Sciences 210 (Agronomy), 230 (Animal Husbandry and Veterinary Science), 310 (Basic Medicine), 320 (Clinical Medicine), 330 (Preventive Medicine and Public Health), 350 (Pharmacy), 360 (Chinese Medicine and Chinese Materia Medica), 413 (Information and System Science Related Engineering and Technology), 416 (Natural Science Related Engineering and Technology), 420 (Surveying and Mapping Science and Technology), 550 (Food Science and Technology)
Table 6: Classification of subjects into STEM (Science, Technology, Engineering, and Mathematics), Social Sciences, Humanities, and Applied Sciences.
A.2 Agreement
Level Agreement ($\kappa\uparrow$)
$m=1$ $m=10$
elementary 0.844 0.838
middle 0.885 0.883
high 0.849 0.846
average 0.864 0.861
Table 7: Cohen’s Kappa agreement ($\kappa$) between GPT-4o and majority voting ($m$: the number of votes) using Qwen2.5-72B-Instruct as evaluator across different education levels of math problems.
Note that for each instance, we have only a single decision from GPT-4o. While it may align more closely with an individual sampled decision from the reward model than with the majority vote (when $m>1$), the latter provides a more stable and deterministic outcome by reducing randomness during grading.
Figure 3: Agreement between GPT-4o and Majority Vote with m Graders, measured by Cohen’s Kappa.
Level Agreement ($\kappa\uparrow$)
$m=1$ $m=10$
college-level 0.881 0.883
Table 8: Cohen’s Kappa agreement ($\kappa$) between GPT-4o and majority voting ($m$: the number of votes) using Qwen2.5-72B-Instruct as evaluator across college-level multi-subject problems.
A.3 REINFORCE
$\nabla_{\theta}\mathbb{E}_{y_{i}\sim\pi_{\theta}(\cdot\mid x)}\Big{[}r(x,a,y_{i})\Big{]}$ $=\sum_{y_{i}}\nabla_{\theta}\Big{[}\pi_{\theta}(y|x)\Big{]}r(x,a,y_{i})$ (7)
$=\sum_{y_{i}}\Big{[}\pi_{\theta}(y|x)\nabla_{\theta}\log\pi_{\theta}(y|x)\Big{]}r(x,a,y_{i})$
$=\mathbb{E}_{y_{i}\sim\pi_{\theta}(\cdot\mid x)}\Big{[}\nabla_{\theta}\log\pi_{\theta}(y_{i}|x)r(x,a,y_{i})\Big{]}.$
A.4 Hyper parameters
Hyperparameter Reward Training Main Experiments
RL SFT RL SFT
micro_train_batch_size 8 4 8 4
train_batch_size 128 128 128 128
micro_rollout_batch_size 16 16
rollout_batch_size 128 128
n_samples_per_prompt 4 4
max_samples 40000 1600000 30000 30000
max_epochs 1 1 1 1
prompt_max_len 1024 1024
generate_max_len 1024 1024
max_len 4096 4096
actor_learning_rate 5e-7 5e-7
init_kl_coef 0.01 0.01
Table 9: Training hyper parameters. Other hyper parameters are the default configuration in OpenRLHF.
Table 9 shows the hyper parameters of our experiments.
coarse fine question answer
Social Sciences Psychology Setting up an activity for students to ’bomb’ each other with compliments belongs to ( ). Self-awareness guidance
STEM Civil Engineering A gravity retaining wall meets the Rankine earth pressure conditions, $H=3\,\mathrm{m}$, top width $2\,\mathrm{m}$, bottom width $3\,\mathrm{m}$, fill $c=0$, $\phi=30^{\circ}$, $\gamma=18.0\,\mathrm{kN/m^{3}}$, the base friction coefficient is 0.4, the anti-sliding stability safety factor $K_{s}$ and the anti-tilting stability safety factor $K_{t}$ are respectively () 2.67; 1.73
Humanities Philosophy Laozi pointed out in the ’Tao Te Ching’, ’Without leaving the door, one knows the world; without peering through the window, one knows the way of heaven. The farther one goes, the less one knows. Therefore, the sage knows without traveling, sees without looking, and achieves without doing.’ Laozi’s view here denies the decisive role of practice in understanding
Applied Sciences Agronomy Under light, the physiological processes that can occur in the mesophyll cells and vascular bundle sheath cells of wheat (C3) are Production of ATP and [H]
Table 10: Example question and reference answer pairs in ExamQA.